OpenStudy (anonymous):

Verify the trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation. (sin x)(tan x cos x - cot x cos x) = 1 - 2 cos^2 x

OpenStudy (anonymous):

tan = sin/cos So tan*cos = (sin/cos)*cos = sin

OpenStudy (anonymous):

Then distribute in sin x.

OpenStudy (anonymous):

so (sin x)(tan x cos x - cot x cos x) (sin x)(sin x - cot x cos x) how would I simplify the cot x cos x part?

hero (hero):

\[\sin(x)\left(\sin(x) - \frac{\cos(x)}{\sin(x)}\cos(x)\right)\]Multiply sin(x) by sinxover sinx: \[\sin(x)\left(\frac{\sin^2(x)}{\sin(x)} - \frac{\cos(x)}{\sin(x)}\cos(x)\right)\]Combine fractions to get:\[\sin(x)\left(\frac{\sin^2(x) - \cos^2x}{\sin(x)} \right)\] Cancel sin x

OpenStudy (anonymous):

how does that equal 1 - 2 cos^2 x ?

OpenStudy (anonymous):

thank you for your help btw

hero (hero):

Did you cancel sin(x) yet?

OpenStudy (anonymous):

I don't think I'm doing it correctly.

hero (hero):

\[\cancel{\sin(x)}\left(\frac{\sin^2(x) - \cos^2x}{\cancel{\sin(x)}} \right)\]

OpenStudy (anonymous):

lol oh i feel stupid. i was trying to cancel out sin^2 x on the numerator.

OpenStudy (anonymous):

I'm still not sure how to get from sin^2 x - cos^2 x to 1-2cos^2 x

hero (hero):

Hint: \(\sin^2x = 1 - \cos^2x\) (by Pythagorean Identity)

OpenStudy (anonymous):

oooh 1 - cos^2 x - cos^2 x = 1- 2 cos^2 x thank you so much